# Octopus-like Reaching Motion: A Perspective Inspired by Whipping

### v1

@@ I INTRODUCTION
The reaching movement of the octopus arm exemplifies one of the most stereotyped and simplified control strategies in its highly deformable muscular system. Unlike the human arm with a limited number of joints, the octopus arm possesses virtually infinite degrees of freedom, functioning as a muscular hydrostat capable of bending, elongating, shortening, and twisting without skeletal support. Despite this structural complexity, octopuses can extend their arms toward a target through highly similar reaching movements, following a consistent bend-propagation pattern that is organized both spatially and temporally [1, 2, 3].
@@
Figure 1: (a) Multiple exposures of a bullwhip exhibiting propagating waves during motion, kinematically similar to an octopus arm. [4] (b) Curvature pattern of real octopus reaching behavior ($s/L_{0}$ denotes the normalized arc length along the arm, with $s$ being the arc-length coordinate and $L_{0}$ the undeformed arm length).
@@
Previous studies have suggested that this movement shares intriguing parallels with whipping-like passive dynamics [1, 2, 5, 6]. Biological reaching can be described as a base-to-tip propagating bend wave—arising from continuous neural drive together with intrinsic elasticity and hydrodynamic interaction—in which a traveling front of stiffening and muscle activation produces curvature and energy transmission [7, 8] reminiscent of a whip crack, as shown in Figure. 1a [4]. The analogy suggests that the octopus may utilize passive mechanical whipping to achieve rapid, efficient reaching movement, converting local activation into a large-scale sweeping-like reaching motion.
@@
The detailed kinematics of the octopus reaching behavior reveal several invariant features across individuals and contexts. The motion is initiated by the formation of a localized bend near the arm base, which then propagates distally along a nearly planar trajectory, maintaining a constant-curvature bend propagation region over a significant portion of the movement, as shown in Figure. 1b [1, 3, 7]. The bend-point velocity shows a characteristic bell-shaped temporal profile with smooth acceleration and deceleration, implying that control is exerted chiefly over the bend’s spatiotemporal position rather than over discrete muscle segments [1]. Biomechanical and computational models further demonstrate that a simple traveling wave of muscle stiffening can reproduce the experimentally observed velocity profiles and curvature evolutions [3, 7], supporting the notion that the system relies on local passive wave mechanics to generate globally coordinated motion. Collectively, these findings portray octopus reaching as a dynamic balance between active control and passive mechanical propagation, embodying principles that resonate strongly with whipping dynamics—where a localized deformation evolves into a smooth, accelerating wave culminating in effective, directed motion.
@@
The whipping motion represents a canonical example of energy focusing and curvature propagation in flexible continua. When an impulse is applied to the handle of a whip, a traveling loop or curvature wave propagates toward the free end, gradually accelerating as the whip’s radius tapers and mass per unit length decreases [9]. This process converts distributed elastic and kinetic energy into a concentrated, high-velocity motion at the tip—eventually reaching supersonic speed and producing the characteristic crack. Theoretical analyses based on elastic rod models reveal that the traveling wave maintains a quasi-constant curvature while local velocity increases, consistent with conservation of energy and momentum in a homogeneous medium. Importantly, this bend propagation mechanism is not unique to whips; similar dynamic principles have been identified in biological systems such as fly-fishing lines, flagella, and even octopus arms, where a local curvature travels through the body to produce directed motion. From this perspective, the octopus reaching behavior may be hypothesis as a biological manifestation of the whipping phenomenon [4]. Both systems rely on the propagation of a localized bend—maintaining nearly constant curvature and exhibiting a bell-shaped velocity evolution. In whips, the traveling loop is a fully passive mechanical wave driven by the initial impulse and geometric tapering; in octopus arms, a similar propagation arises from the coupling of muscle activation waves and passive environmental interactions. These parallels suggest that understanding whipping dynamics provides not only a physical analogue but also an experimental framework for interpreting the passive wave-based control strategy observed in octopus reaching motions.
@@ II-B Data Extraction
Midline Extraction. To quantitatively analyze the deformation of the silicone octopus arm, we developed a custom image-processing pipeline. As shown in Figure. 4a, each video frame was first processed with Meta’s Segment Anything Model 2 (SAM 2) [10] to segment the moving arm from the underwater background. A binary mask was then applied to remove the rigid supporting structure, yielding a clean extraction of the silicone arm (Figure. 4b). The extracted region was skeletonized to obtain its medial axis (Figure. 4c). Because skeletonization can occasionally introduce spurious branches, all branch points were detected and removed. The three terminal nodes of the remaining skeleton were subsequently identified, and the two closest endpoints $(E_{2},E_{3})$ were used to compute their midpoint $M$, which served as an auxiliary node to reconstruct the main, ordered centerline (Figure. 4d).
@@ III RESULTS
(a) Normalized velocity pattern for velocity and distance and aligned at the peaks.[1]
@@ IV DISCUSSION
Interestingly, the curvature patterns of Gel 2 at low driving speed (30 rpm) exhibited temporal profiles similar to those of Ecoflex 10 and Dragonskin 10 driven at 150 rpm. This correspondence suggests that curvature propagation depends not only on material modulus but also on the relative balance between actuation intensity and inertial loading. The softer Gel 2 at low driving speed and the stiffer Ecoflex 10 and Dragonskin 10 at moderate speed appear to reach comparable dynamic regimes where elastic recovery dominates over inertia-driven momentum transfer. Under these conditions, the bending wave travels through passive redistribution of kinetic and elastic energy along the arm, leading to similar curvature attenuation profiles across materials. This behavior aligns with the whip-wave dynamics described by McMillen and Goriely [9], where energy imparted at the base propagates through a continuous medium via inertial coupling and geometric nonlinearity.
@@
Furthermore, at higher speeds (210 rpm), the silicone arm occasionally exhibited non-planar motion characterized by twisting and asymmetric trajectories. We hypothesize that this phenomenon arises from strong fluid–structure coupling, where the interaction between the rapidly deforming arm and the surrounding flow generates three-dimensional hydrodynamic instabilities. Similar out-of-plane deformations have been reported in previous fluid–structure interaction studies of flexible bodies in oscillatory or traveling-wave motion [11, 12]. In these works, the coupling between structural elasticity, inertial loading, and unsteady vortex shedding can drive spontaneous transitions from planar to helical or chaotic motion once the driving velocity exceeds a critical threshold. In our system, this transition likely reflects the onset of such coupling-induced instability, suggesting that fluid feedback becomes dynamically significant beyond the regime of quasi-planar propagation.
@@
These findings indicate that the bend propagation observed in our experiments can be parsimoniously explained by a passive hydrodynamic mechanism. Once a local bend is initiated near the base, distributed drag from the surrounding water effectively suppresses the unactuated portions of the arm and advects the bend distally along a nearly planar path. The absence of propagation in air under identical conditions (Gel 2 driven at 150 rpm), as shown in Figure. 8a, further supports this interpretation. In addition, experiments in the attached video conducted in air with driving speeds at 210, 270, and 330 rpm (see Figure. 8b) revealed bend propagation patterns that differed substantially from those observed in biological reaching, accompanied by a marked increase in bend-point velocity [9]. This suggests that, in the absence of fluid-mediated damping and added-mass effects, the motion is dominated by direct mechanical inertia rather than hydrodynamic coupling. Therefore, we speculate that the surrounding fluid plays an important role in shaping curvature propagation, consistent with previous biological findings describing passive wave transmission in “non-activated” octopus arm segments [3].

### v2

@@ I INTRODUCTION
The reaching movement of the octopus arm exemplifies one of the most stereotyped and simplified control strategies in its highly deformable muscular system. Unlike the human arm with a limited number of joints, the octopus arm possesses virtually infinite degrees of freedom, functioning as a muscular hydrostat capable of bending, elongating, shortening, and twisting without skeletal support. Despite this structural complexity, octopuses can extend their arms toward a target through highly similar reaching movements, following a consistent bend-propagation pattern that is organized both spatially and temporally [gutfreund1996organization, sumbre2001control, yekutieli2005dynamic].
@@
Figure 1: (a) Multiple exposures of a bullwhip exhibiting propagating waves during motion, kinematically similar to an octopus arm. [noel2018grip] (b) Curvature pattern of real octopus reaching behavior ($s/L_{0}$ denotes the normalized arc length along the arm, with $s$ being the arc-length coordinate and $L_{0}$ the undeformed arm length).
@@
Previous studies have suggested that this movement shares intriguing parallels with whipping-like passive dynamics [gutfreund1996organization, sumbre2001control, gutfreund1998patterns, alupay2023syntactic]. Biological reaching can be described as a base-to-tip propagating bend wave—arising from continuous neural drive together with intrinsic elasticity and hydrodynamic interaction—in which a traveling front of stiffening and muscle activation produces curvature and energy transmission [zelman2013kinematic, wang2022control] reminiscent of a whip crack, as shown in Figure. 1a [noel2018grip]. The analogy suggests that the octopus may utilize passive mechanical whipping to achieve rapid, efficient reaching movement, converting local activation into a large-scale sweeping-like reaching motion.
@@
The detailed kinematics of the octopus reaching behavior reveal several invariant features across individuals and contexts. The motion is initiated by the formation of a localized bend near the arm base, which then propagates distally along a nearly planar trajectory, maintaining a constant-curvature bend propagation region over a significant portion of the movement, as shown in Figure. 1b [gutfreund1996organization, yekutieli2005dynamic, zelman2013kinematic]. The bend-point velocity shows a characteristic bell-shaped temporal profile with smooth acceleration and deceleration, implying that control is exerted chiefly over the bend’s spatiotemporal position rather than over discrete muscle segments [gutfreund1996organization]. Biomechanical and computational models further demonstrate that a simple traveling wave of muscle stiffening can reproduce the experimentally observed velocity profiles and curvature evolutions [yekutieli2005dynamic, zelman2013kinematic], supporting the notion that the system relies on local passive wave mechanics to generate globally coordinated motion. Collectively, these findings portray octopus reaching as a dynamic balance between active control and passive mechanical propagation, embodying principles that resonate strongly with whipping dynamics—where a localized deformation evolves into a smooth, accelerating wave culminating in effective, directed motion.
@@
The whipping motion represents a canonical example of energy focusing and curvature propagation in flexible continua. When an impulse is applied to the handle of a whip, a traveling loop or curvature wave propagates toward the free end, gradually accelerating as the whip’s radius tapers and mass per unit length decreases [mcmillen2003whip]. This process converts distributed elastic and kinetic energy into a concentrated, high-velocity motion at the tip—eventually reaching supersonic speed and producing the characteristic crack. Theoretical analyses based on elastic rod models reveal that the traveling wave maintains a quasi-constant curvature while local velocity increases, consistent with conservation of energy and momentum in a homogeneous medium. Importantly, this bend propagation mechanism is not unique to whips; similar dynamic principles have been identified in biological systems such as fly-fishing lines, flagella, and even octopus arms, where a local curvature travels through the body to produce directed motion. From this perspective, the octopus reaching behavior may be hypothesis as a biological manifestation of the whipping phenomenon [noel2018grip]. Both systems rely on the propagation of a localized bend—maintaining nearly constant curvature and exhibiting a bell-shaped velocity evolution. In whips, the traveling loop is a fully passive mechanical wave driven by the initial impulse and geometric tapering; in octopus arms, a similar propagation arises from the coupling of muscle activation waves and passive environmental interactions. These parallels suggest that understanding whipping dynamics provides not only a physical analogue but also an experimental framework for interpreting the passive wave-based control strategy observed in octopus reaching motions.
@@ II-B Data Extraction
Midline Extraction. To quantitatively analyze the deformation of the silicone octopus arm, we developed a custom image-processing pipeline. As shown in Figure. 4a, each video frame was first processed with Meta’s Segment Anything Model 2 (SAM 2) [ravi2024sam2] to segment the moving arm from the underwater background. A binary mask was then applied to remove the rigid supporting structure, yielding a clean extraction of the silicone arm (Figure. 4b). The extracted region was skeletonized to obtain its medial axis (Figure. 4c). Because skeletonization can occasionally introduce spurious branches, all branch points were detected and removed. The three terminal nodes of the remaining skeleton were subsequently identified, and the two closest endpoints $(E_{2},E_{3})$ were used to compute their midpoint $M$, which served as an auxiliary node to reconstruct the main, ordered centerline (Figure. 4d).
@@ III RESULTS
(a) Normalized velocity pattern for velocity and distance and aligned at the peaks.[gutfreund1996organization]
@@ IV DISCUSSION
Interestingly, the curvature patterns of Gel 2 at low driving speed (30 rpm) exhibited temporal profiles similar to those of Ecoflex 10 and Dragonskin 10 driven at 150 rpm. This correspondence suggests that curvature propagation depends not only on material modulus but also on the relative balance between actuation intensity and inertial loading. The softer Gel 2 at low driving speed and the stiffer Ecoflex 10 and Dragonskin 10 at moderate speed appear to reach comparable dynamic regimes where elastic recovery dominates over inertia-driven momentum transfer. Under these conditions, the bending wave travels through passive redistribution of kinetic and elastic energy along the arm, leading to similar curvature attenuation profiles across materials. This behavior aligns with the whip-wave dynamics described by McMillen and Goriely [mcmillen2003whip], where energy imparted at the base propagates through a continuous medium via inertial coupling and geometric nonlinearity.
@@
Furthermore, at higher speeds (210 rpm), the silicone arm occasionally exhibited non-planar motion characterized by twisting and asymmetric trajectories. We hypothesize that this phenomenon arises from strong fluid–structure coupling, where the interaction between the rapidly deforming arm and the surrounding flow generates three-dimensional hydrodynamic instabilities. Similar out-of-plane deformations have been reported in previous fluid–structure interaction studies of flexible bodies in oscillatory or traveling-wave motion [wang2020review, ma2022flexible]. In these works, the coupling between structural elasticity, inertial loading, and unsteady vortex shedding can drive spontaneous transitions from planar to helical or chaotic motion once the driving velocity exceeds a critical threshold. In our system, this transition likely reflects the onset of such coupling-induced instability, suggesting that fluid feedback becomes dynamically significant beyond the regime of quasi-planar propagation.
@@
These findings indicate that the bend propagation observed in our experiments can be parsimoniously explained by a passive hydrodynamic mechanism. Once a local bend is initiated near the base, distributed drag from the surrounding water effectively suppresses the unactuated portions of the arm and advects the bend distally along a nearly planar path. The absence of propagation in air under identical conditions (Gel 2 driven at 150 rpm), as shown in Figure. 8a, further supports this interpretation. In addition, experiments in the attached video conducted in air with driving speeds at 210, 270, and 330 rpm (see Figure. 8b) revealed bend propagation patterns that differed substantially from those observed in biological reaching, accompanied by a marked increase in bend-point velocity [mcmillen2003whip]. This suggests that, in the absence of fluid-mediated damping and added-mass effects, the motion is dominated by direct mechanical inertia rather than hydrodynamic coupling. Therefore, we speculate that the surrounding fluid plays an important role in shaping curvature propagation, consistent with previous biological findings describing passive wave transmission in “non-activated” octopus arm segments [yekutieli2005dynamic].